Knowing that P = 120 lb and that the resultant of P and Q lies in the positive x-direction,

determine Q and the magnitude of the resultant.


1
Expert's answer
2021-09-08T16:13:44-0400


From Force body diagram

"sinθ = \frac{3}{5} \\

cosθ = \frac{4}{5} \\

sinϕ = \frac{5}{13} \\

cosϕ = \frac{12}{13} \\

P= 120 \;lb"

Since there is resultant only in x-direction and no resultant force in y-direction.

"P_y=Q_y \\

Psinθ =Qcosϕ \\

120 \times \frac{3}{5}= Q \times \frac{12}{13} \\

Q = 120 \times \frac{3}{5} \times \frac{13}{12} \\

Q = 78 \;lb"

Resultant

"R=Q_x+P_x \\

R= Qsinϕ + Pcosθ \\

R= (78 \times \frac{5}{13}) + (120 \times \frac{4}{5}) \\

= 30+96 = 126 \;lb"